OpenStudy (anonymous):

.'.'. i will give medal.'.'. 4x-3y=14 2x+y=-3 i need whole solution.

OpenStudy (anonymous):

yes?

OpenStudy (juanpablojr):

medal?

OpenStudy (anonymous):

what is the ?

OpenStudy (anonymous):

look at the ques

OpenStudy (juanpablojr):

4x+3y=1, 3x+4y=10 - x=-26/7, y = 37/7

OpenStudy (anonymous):

you could use substitution so isolate y in the second ? then plug in what you get after that into your first y then solve from there so it would be 4x+3(-3-2x)=14 after that you distribute so it becomes 4x-6x=14+9. which leads to -2x=23)=x=-11.5 after that plug x into one of the equations to get y so I pick the 1st equation so it becomes 4(-11.5)+3y=14 you solve for y so it becomes 3y=60 y=20